# Pressure induced redistribution of oxygen hole states in La4Ni3O10

### 14 Mar 2026 v1

@@ Abstract
Using density functional calculations and multi-orbital, multi-atom cluster exact diagonalization that includes local exchange and Coulomb interactions, we explored the local low-energy electronic states of trilayer La4Ni3O10 via a minimal Ni3O14 cluster. We find that, at ambient pressure, starting with all three Ni being nominally 2+ valence, one of the two extra holes is localized in the central NiO2 layer forming a Zhang-Rice singlet (ZRS) with $d_{x^{2}-y^{2}}$ orbital. The other hole mainly occupies the antibonding combination of the two interplane O $p_{z}$ orbitals and thereby hybridizes with an out-of-plane three-spin-polaron (3SP) formed by the $d_{z^{2}}$ orbitals of three NiO2 layers. In this way, the in-plane spin orientation alternation is carried by the $d_{x^{2}-y^{2}}$ orbitals of two outer layers with interlayer antiferromagnetic correlation. Simultaneously, the central layer is insulating like with negligible magnetic moment. At high pressure, however, the two extra holes are concentrated on one of two outer layers and the inner layer separately forming the ZRS with $d_{x^{2}-y^{2}}$ orbitals or in-plane 3SP with neighboring cluster. We highlight the similarities between the bilayer La3Ni2O7 and trilayer La4Ni3O10 via possible charge and spin ordered states suggested by our cluster results.
@@ II.1 DFT calculations and Wannier projections
Density-functional theory  [9, 14] level structural relaxations and band structure calculations were performed using the VASP  [16, 15] and Wien2K  [1, 25]codes with the Perdew-Burke-Ernzerhof version of the generalized gradient approximation (GGA-PBE)  [24] and a dense of over 2000 $k$-points for the unit cell of La4Ni3O10. The parameters listed in Table I are determined from the corresponding DFT Bloch bands and Maximally localized Wannier functions  [31, 21] projections using Wannier90  [23] and WIEN2WANNIER  [17] interface. The most crucial parameters in all three models are the hybridization tds between $d_{x^{2}-y^{2}}, Ni-3$$_{dz^{2}}$ and in-plane and apical O-2pp orbitals, and these hopping parameters between Ni-3dd orbitals, Ni-3dd/O-2pp to O-2pp were obtained via the Wannier projections for models of Ni-3dd subspace and Ni-3dd and O-2pp substate, respectively.
@@ II.2 Multi-orbital cluster Model
Assuming a fixed 2-2 valence for oxygen, the total formal valence of the three Ni ions in La4Ni3O10 sums up to +8+8, producing an average Ni valence of +2.67+2.67, commonly described as a mixture of Ni2+ and Ni3+. However, compared to the bilayer La3Ni2O7, the hole distribution in this trilayer system is considerably more complex. Presumably, in the hole language as adopted in our previous work [11], 6 of 8 holes would occupy the three Ni $3d^{8}$ sites so that the key question is where the remaining two holes locate among different layers and orbitals, which is largely affected by the pressure induced change of various hybridization.
@@ II.3 Energy level diagram
$E(d^{6})$ $=E(d^{10})=2U.$ (8)
Figure 2: Energy level diagram of 8-hole cluster in the absence of hybridization between configurations. The red (blue) arrow indicates $t_{dO}$ ($t_{pd}$) hybridization. The energy of $d^{8}$ is chosen as the zero energy reference.
@@ III Results and discussion
In addition, the left side of Fig. 3 also reveals the in-plane spin orientation alternation carried by the two outer layers with interlayer antiferromagnetic correlation. Concurrently, the central layer is insulating similar to what happens in the 113 nickelates like NdNiO3 (add reference). These observations are all consistent with the recent experiments [34, 36, 8].

### 18 Mar 2026 v2

@@ Abstract
Using density functional calculations and multi-orbital, multi-atom cluster exact diagonalization that includes local exchange and Coulomb interactions, we explored the local low-energy electronic states of trilayer La4Ni3O10 via a minimal Ni3O14 cluster. We find that, at ambient pressure, starting with all three Ni being nominally 2+ valence, one of the two extra holes is localized in the central NiO2 layer forming a Zhang-Rice singlet (ZRS) with $d_{x^{2}-y^{2}}$ orbital. The other hole mainly occupies the antibonding combination of the two interplane O $p_{z}$ orbitals and thereby hybridizes with an out-of-plane three-spin-polaron (3SP) formed by the $d_{z^{2}}$ orbitals of three NiO2 layers. In this way, the in-plane spin orientation alternation is carried by the $d_{x^{2}-y^{2}}$ orbitals of two outer layers with interlayer antiferromagnetic correlation. Simultaneously, the central layer is insulating with negligible magnetic moment. At high pressure, however, the two extra holes are concentrated on one of two outer layers and the inner layer separately forming the ZRS with $d_{x^{2}-y^{2}}$ orbitals or in-plane 3SP with neighboring cluster. We highlight the similarities between the bilayer La3Ni2O7 and trilayer La4Ni3O10 via possible charge and spin ordered states suggested by our cluster results.
@@ II.1 DFT calculations and Wannier projections
The Density-functional theory [9, 14] level structural relaxations and band structure calculations were performed using the VASP [16, 15] and Wien2K [1, 25] codes with the Perdew-Burke-Ernzerhof version of the generalized gradient approximation (GGA-PBE) [24] and a dense of over 2000 $k$-points for the unit cell of La4Ni3O10. The parameters listed in Table I are determined from the corresponding DFT Bloch bands and Maximally localized Wannier functions [31, 21] projections using Wannier90 [23] and WIEN2WANNIER [17] interface. The most crucial parameters in all three models are the hybridization tds between $d_{x^{2}-y^{2}}, Ni-3$$_{dz^{2}}$ and in-plane and apical O-2pp orbitals, and these hopping parameters between Ni-3dd orbitals, Ni-3dd/O-2pp to O-2pp were obtained via the Wannier projections for models of Ni-3dd subspace and Ni-3dd and O-2pp substate, respectively.
@@ II.2 Multi-orbital cluster Model
Assuming a fixed 2-2 valence for oxygen, the total formal valence of the three Ni ions in La4Ni3O10 sums up to +8+8, producing an average Ni valence of +2.67+2.67, commonly described as a mixture of Ni2+ and Ni3+. However, compared to the bilayer La3Ni2O7, the hole distribution in this trilayer system is considerably more complex. Presumably, in the hole language adopted in our previous work [11], 6 of 8 holes would occupy the three Ni $3d^{8}$ sites so that the key question is where the remaining two holes locate among different layers and orbitals, which is largely affected by the pressure induced change of various hybridization.
@@ II.3 Energy level diagram
$E(d^{6})$ $=E(d^{10})=2U$ (8)
@@
Figure 2: Energy level diagram of 8-hole cluster in the absence of hybridization and correlations apart from the Ni $U$. The red (blue) arrow indicates $t_{dO}$ ($t_{pd}$) hybridization matrix elements. The energy of $d^{8}$ is chosen as the zero energy reference. These configurations are highly degenerate and this degenerate will be lifted when switching on the Hund’s rule exchange and various hopping integrals. Configurations with dominant weights at DFT value in the 8-hole are highlighted in red. Configurations with even higher energies are not shown for clarity.
@@ III Results and discussion
In addition, the left side of Fig. 3 also reveals the in-plane spin orientation alternation carried by the two outer layers with interlayer antiferromagnetic correlation. Concurrently, the central layer is insulating similar to what happens in the 113 nickelates like NdNiO3 [22]. These observations are all consistent with the recent experiments [34, 36, 8].